OpenStudy (anonymous):

If f(x) = x + 4 and g(x) = 2x - 3, find (g - f)(2). Write the answer as an integer.

OpenStudy (michele_laino):

by definition we can write: \[\left( {g - f} \right)\left( x \right) = g\left( x \right) - f\left( x \right) = \left( {2x - 3} \right) - \left( {x + 4} \right) = ...?\]

OpenStudy (anonymous):

okay

OpenStudy (michele_laino):

hint: \[\begin{gathered} \left( {g - f} \right)\left( x \right) = g\left( x \right) - f\left( x \right) = \left( {2x - 3} \right) - \left( {x + 4} \right) = \hfill \\ \hfill \\ = 2x - 3 - x - 4 = ...? \hfill \\ \end{gathered} \]

OpenStudy (anonymous):

I got 6

OpenStudy (michele_laino):

I got a different result

OpenStudy (anonymous):

I got 9

OpenStudy (michele_laino):

hint: \[\begin{gathered} \left( {g - f} \right)\left( x \right) = g\left( x \right) - f\left( x \right) = \left( {2x - 3} \right) - \left( {x + 4} \right) = \hfill \\ \hfill \\ = 2x - 3 - x - 4 = x - 7 \hfill \\ \end{gathered} \]

OpenStudy (michele_laino):

now, replace x with 2, please what do you get?

OpenStudy (anonymous):

-5

OpenStudy (michele_laino):

that's right!

OpenStudy (anonymous):

so would the answer be -5

OpenStudy (michele_laino):

yes!

OpenStudy (anonymous):

If f\left( x \right) = \frac{{2{x^2}}}{{x - 1}}, evaluate for f(-4). Write only the answer as a decimal or fraction. also got stuck on this question

OpenStudy (michele_laino):

is your function, like this: \[f\left( x \right) = \frac{{2{x^2}}}{{x - 1}}\]

OpenStudy (anonymous):

yes

OpenStudy (michele_laino):

we have to replace x with -4, so we can write: \[f\left( { - 4} \right) = \frac{{2{{\left( { - 4} \right)}^2}}}{{ - 4 - 1}} = ...?\]

OpenStudy (anonymous):

-13

OpenStudy (michele_laino):

hint: \[f\left( { - 4} \right) = \frac{{2{{\left( { - 4} \right)}^2}}}{{ - 4 - 1}} = \frac{{2 \times 16}}{{ - 5}} = ...?\]

OpenStudy (anonymous):

27

OpenStudy (michele_laino):

are you sure? what is 2*16=...?

OpenStudy (anonymous):

32

OpenStudy (michele_laino):

correct! so our answer is: \[f\left( { - 4} \right) = \frac{{2{{\left( { - 4} \right)}^2}}}{{ - 4 - 1}} = \frac{{2 \times 16}}{{ - 5}} = \frac{{32}}{{ - 5}} = - \frac{{32}}{5}\]

OpenStudy (anonymous):

okay